1. Write the multisets of prime factors for the given numbers.
I. 160
II. 120
III. 250
2. Write the multiplicities of each element of multisets in part 2(1-I, ii,iii) separately.
3. Find the cardinalities of each multiset in part 2-1.

2. Function converts Fahrenheit temperatures into Celsius. What is the function for opposite conversion?
1
Expert's answer
2020-06-09T19:09:06-0400

1.

"160 : \{2,2,2,2,2,5\}"

multiplicity of 2 is 5, multiplicity of 5 is 1

Cardinality"=6"


"120 : \{2,2,2,3,5\}"

multiplicity of 2 is 3, multiplicity of 3 is 1, multiplicity of 5 is 1

Cardinality"=5"


"250 : \{2,5,5,5\}"

multiplicity of 2 is 1, multiplicity of 5 is 3

Cardinality"=4"


2.

Function converts Fahrenheit temperatures into Celsius:

"T_{\degree C}=\frac {5}{9}(T_{(\degree F)}-32)"

Then:

"T_{(\degree F)}-32=\frac {9}{5}T_{(\degree C)}"

The function for opposite conversion:

"T_{(\degree F)}=\frac {9}{5}T_{(\degree C)}+32"